what is the result of multiplying 100 more than the difference between 4321 and 3194 by 11
Answer by Theo(13342)

the difference between 4321 and 3194 is equal to 4321 - 3194 = 1127.

100 more than this would be 1127 + 100 = 1227.

11 * 1227 = 13,497.

if you wanted to add parentheses to the words, you would probably do the following:

what is the result of multiplying (100 more than the (difference between 4321 and 3194)) by 11.

first you find the difference.
then you add 100 to it.
then you multiply the result.

if you look at the word multiplying and then the word by, it's the stuff in between that's being multiplied.

if you look at the words more than the difference, you would find the difference first and then add 100 to it.

that's my take.

this one is fairly straight forward in the translations.

others can be a bit more murky.